OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 4x + 2y = 6 x – y = 3 −1 5 −2 2

OpenStudy (anonymous):

well, where r u getting stuck?

OpenStudy (anonymous):

The whole thing

OpenStudy (anonymous):

lol okki

OpenStudy (anonymous):

4x + 2y = 6 x – y = 3 4x + 2y = 6 2x – 2y = 6 clear till here?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

now add them both

OpenStudy (anonymous):

all of it

OpenStudy (anonymous):

4x + 2y = 6 and, 2x – 2y = 6

OpenStudy (anonymous):

ok 6x+4y=12

OpenStudy (anonymous):

nahh, +2y - 2y = ??

OpenStudy (anonymous):

this is a simaltenious equation x=2.

OpenStudy (anonymous):

ok whats next

OpenStudy (anonymous):

y will get cancelled, did u get that??

OpenStudy (anonymous):

no

OpenStudy (anonymous):

you substitute x with 2 and u will get your y

OpenStudy (anonymous):

so the answer is 2?

OpenStudy (anonymous):

4x + 2y = 6 + 2x – 2y = 6 4x + 2x + 2y - 2y = 12 6x = 12 x = 2

OpenStudy (anonymous):

Thanks to the both of you

OpenStudy (anonymous):

you're welcome

OpenStudy (anonymous):

anytime sammy